OpenStudy (anonymous):

write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5

OpenStudy (anonymous):

x-intercept (6,0) y-intercept (0,5) now find the gradient from those two points and substitute a point from above and the gradient back into the formula y=mx+c so find the value of c and then rewrite the whole equation with the c value that you found.

OpenStudy (anonymous):

I like to break these down. So slope intercept form is: y = mx + b And m is the slope and b is the y-intercept They give us the y intercept, so we can start filling in the equation: y = mx + 5 So now we substitute in values for x and y to solve for m: We know the x-intercept is 6, so that's the point (6,0), where x=6 and y=0, so let's substitute those into our equation so far: 0 = m(6) + 5 -5 = 6m m = -5/6 So the equation is: y = (-5/6)x + 5 Hope that helps! :-) It's my first day on the site.

OpenStudy (anonymous):

Thanks guys